OpenStudy (anonymous):

Harold has three dogs. The Irish terrier weighs one-third as much as the golden retriever. The golden retriever weighs half as much as the Saint Bernard. If Harold's Saint Bernard weighs one hundred fifty pounds, what is the combined weight of all three dogs? two hundred lb two hundred twenty-five lb two hundred fifty lb two hundred seventy-five lb

OpenStudy (anonymous):

Two hundred fifty lb Because golden retriever = half as much as SaintB = 150/2=75lb Irish terrier = 1/3 of golden retriever = 75/3 =25 Therefore weight of SaintB+IrishT+GoldenR = 150+75+25= 250lb.